# Artifact-free Sound Quality in DNN-based Closed-loop Systems for Audio Processing

### v2

@@ III-B Auditory modules and HA model
When selecting hyperparameters, two primary considerations should be addressed: the receptive field (RF) and the specific activation function. A larger receptive field (RF) ensures that the CNN-based network accurately simulates the auditory model’s impulse response. The selected activation functions need to capture the nonlinear characteristics of auditory processing, such as the compressive growth of the BM vibrations in the cochlear model and the rectification-like behavior of the IHCs. The framework’s hyperparameters are detailed in Table I. Specifically, $K_{1}$ and $K_{2}$ denote the kernel sizes of the dilated depthwise convolution layers (Dh-conv and Df-conv) within each memory block. The parameter $H$ indicates the number of channels in the hidden layers. Additionally, $L$, $L_{l}$, and $L_{r}$ represent the input length, left context, and right context, respectively.
The dCoNNearcochlear transforms acoustic signals into cochlear basilar-membrane vibrations. As described in [20], the hyperbolic tangent (Tanh) activation function was used, because it allows the activations to resemble the input-output relations of the auditory model and supports both positive and negative deflections of the basilar membrane.
The dCoNNearIHC model emulates the IHCs’ function of sensing basilar membrane vibrations and converting them into IHC receptor potential changes. It employs a Tanh nonlinearity in the first stack and a sigmoid function in the second stack. This approach mirrors the compressive characteristics of the IHC input/output functions, with the sigmoid function effectively capturing the IHC receptor potential as a negative voltage difference [6].
The dCoNNearANF model converts IHC receptor potentials into firing rates for three types of ANFs. The initial M blocks are shared layers, with each branch containing M blocks to independently generate HSR, MSR, and LSR fiber responses. The overall AN response ($r_{f}$) is derived by summing the three ANF responses. The ReLU activation function is applied here, as ANF processing differs from cochlear and IHC processing in that it only involves positive output values, without the compressive properties observed in the previous stages. The model features substantial left context windows of 7936 samples, ensuring accurate capture of the time-dependent recovery properties of the ANF models [26].
@@
TABLE I: Hyperparamters of the dCoNNears and HA-model. M and R represent the number of memory blocks and the repetition number; K1 and K2 denote the kernel size of Dh-conv and Df-conv in each memory block; H indicates the channel number of hidden layers; L, Ll, and Lr represent the input length, left context, and right context, respectively
@@ III-C Individualisation of hearing impairment
To individualise hearing impaired pathway as shown in Fig. 2, we adjusted the normal hearing (NH) auditory modules to simulate different degrees of outer hair cell (OHC) loss and cochlear synaptopathy (CS) in the HI pathway. To simulate OHC loss, we retrained the NH cochlear model using transfer learning, based on a specific gain-loss profile or an individual audiogram [27]. The NH auditory nerve (AN) response ($r_{f}$) is computed as a weighted sum of the three ANF responses ($H_{NH}$ = 13, $M_{NH}$ = 3, and $L_{NH}$ = 3). To simulate CS, the model was modified with adjusted weights $H_{HI}$, $M_{HI}$, and $L_{HI}$ [6]. These personalized adjustments ensured the models accurately represented individual hearing impairments, providing tailored audio processing solutions that minimized artifacts and preserved fidelity.
@@ IV Training strategy
The training procedure involved two main stages: (1) training the auditory modules (dCoNNearcochlear, dCoNNearIHC, and dCoNNearANF), and (2) training the HA model based on the closed-loop framework including the NH and HI pathways. Firstly, we trained the auditory modules using simulated responses from analytical models as targets, following the procedure outlined in [20, 26]. Once the auditory elements were trained, their parameters were frozen, and the HA models were then optimized within the closed-loop system as first described in [6].
@@ IV-A Training auditory elements
The auditory module training targets were derived from various analytical models: a biophysical transmission line (TL) model for dCoNNearcochlear [22], an analytical Hodgkin–Huxley-type model for dCoNNearIHC [28], and a three-store diffusion model of the ANF synapse for dCoNNearANF [29]. To train these models, we used 2,310 utterances from the TIMIT speech corpus [30]. The input signals were first upsampled to 100 kHz to accurately solve the reference models, and then downsampled to 20 kHz for training. The root-mean-square (RMS) energy of each sentence was adjusted to a specified sound pressure level (SPL) using the following equation:
@@ IV-A1 Training dCoNNearcochlear
The dCoNNearcochlear model was trained to simulate basilar membrane (BM) displacements for 201 center frequencies between 112 Hz and 12 kHz [22], as human hearing sensitivity diminishes above 12 kHz. The RMS energy of each sentence was adjusted to $L_{D}=70$ dB SPL. The input signal and TL-model outputs were segmented into windows of 2,048 samples, with left ($L_{l}$) and right ($L_{r}$) contexts of 256 samples each, resulting in a total input length of 2,560 samples. To facilitate training, a scaling factor of $10^{6}$ was applied to the simulated BM displacements (expressed in [$\mu m$]), ensuring that the datasets maintained a statistical mean near zero and a standard deviation close to one.
@@ IV-A2 Training dCoNNearIHC and dCoNNearANF
[…]
@@ IV-C Loss function
The training strategies from [20] and [26] were employed to train the dCoNNearcochlear, dCoNNearIHC, and dCoNNearANF models. The loss function for these models was optimized using the mean absolute error (MAE) between the analytical outputs and the predicted dCoNNear outputs.
For training the HA model parameters, we utilized a combined loss function that accounted for different representations of the differences between the NH responses $r_{f}$ and the HI $\hat{r_{f}}$. As outlined in [6], additional constraints, such as population responses, were introduced to the training process to minimize features functionally relevant to hearing loss and auditory perception, given the highly nonlinear nature of auditory processing. The combined loss function incorporated both the AN responses and the AN population responses across all simulated CFs:
@@ IV-D Hearing impaired elements
We optimized the HA model to compensate for a specific hearing-impaired (HI) profile “Slope35-7,0,0” [6]. This profile represents a sloping high-frequency outer hair cell (OHC) loss beginning at 1 kHz, with hearing thresholds of 35 dB HL at 8 kHz, referred to as Slope35. Additionally, there is a complete loss of low spontaneous rate (LSR) and medium spontaneous rate (MSR) auditory nerve fibers (ANFs) (HHI = 0 and MHI= 0) and approximately 46% loss of high spontaneous rate (HSR) ANFs (LHI = 7) to simulate the effects of age-related cochlear synaptopathy.
@@ IV-E TrainingSetup
The learning rates for all model training was set to $10^{-4}$. If the validation loss did not decrease over five consecutive epochs, the learning rate was halved to ensure optimal training progression. All models were trained for 50 epochs using the Adam optimizer [31]. Training was conducted on three NVIDIA A30 GPUs using the PyTorch platform.
@@ V-A Evaluating Artifacts
To determine whether the artifacts caused by the prior CoNNear were eliminated, we analyzed 70-dB SPL 1-kHz tone responses at various auditory processing stages, as well as the HA models trained within the closed-loop system based on both prior CoNNear and dCoNNear architectures. We quantified artifacts using total harmonic distortion (THD) at each step of the closed-loop system [32]. As aliased harmonics are generated by the prior CoNNear models in the 1-kHz tone responses as shown in Fig. 3, we examined harmonics at one-quarter intervals of the fundamental frequency:
@@
The Speech-to-Reverberation Modulation Energy Ratio (SRMR) [33] was used to evaluate the sound quality of the resulting HA models trained using both the prior CoNNear and dCoNNear-based frameworks. SRMR is a non-intrusive metric that assesses speech quality and intelligibility based on a modulation spectral representation of the speech signals. Higher SRMR scores indicate higher sound quality and fewer artifacts. We evaluated sound quality across three audio types: clean speech, noisy speech, and music. The clean speech dataset was 550 randomly selected utterances from TIMIT. The one-hour noisy speech dataset was generated by mixing the clean speech samples from TIMIT and the noise from Freesound [34] at SNR of 0 dB. The music dataset consisted of 100 randomly selected 30-second tracks from FMA [35].
@@ V-B Evaluating Auditory Elements
For the dCoNNearcochlear, we simulated the QERB and basilar membrane (BM) excitation patterns, as described in [20]. The QERB metric characterizes the level-dependent cochlear filter properties derived from the BM impulse response and is calculated as follows:
@@
For the dCoNNearANF, rate-level curves and synchrony levels were measured [6]. Rate-level curves evaluate ANF responses to changes in stimulus level. We generated pure-tone stimuli (50-ms duration, 2.5-ms rise/fall ramp) with levels between 0 and 100 dB at frequencies of approximately 1 and 4 kHz, based on the corresponding CFs of the ANF models (1007 and 3972.7 Hz). The rate-level functions were derived by computing the average response 10–40 ms after stimulus onset. The ANF synchrony level describes the non-monotonic relation between ANF response and the stimulus level. Fully modulated 400-ms long pure tones with a modulation frequency $f_{m}$ of 100 Hz and carrier frequencies of 3972.7 Hz (henceforth referred to as 4 kHz) were simulated, and the synchrony-level functions were calculated by extracting the magnitude of the $f_{m}$ component from the Fourier spectrum of the fibers’ firing rate.
@@ V-C Evaluating HA models
To evaluate the restoration performance of the trained HA models, we used normalized root-mean-square error (NRMSE) in [6], which was computed between the simulated NH and HI AN population responses, normalised to the maximum of the NH response for each sentence:
@@
NRMSE calculations were performed on 550 randomly selected clean speech utterances from TIMIT, with RMS energy levels ranging from 40 to 70 dB SPL in 10 dB increments.
@@ VI Results
In this section, we evaluated the proposed artifact-free dCoNNear-based closed-loop framework, designed for individualized DNN-based hearing aid (HA) models. The evaluation aimed to determine whether the artifacts associated with the prior CoNNear architecture were successfully eliminated in the dCoNNear-based auditory elements. We also compared the sound quality of the HA models trained with both the prior CoNNear and the new dCoNNear frameworks. Additionally, we assessed the biophysical properties of the auditory elements using stimuli that were not part of the training material. Finally, we compared the hearing aid restoration performance of the models developed using the prior CoNNear-based framework with those developed using the dCoNNear framework.
@@ VI-A Artifacts and sound quality
Table LABEL:tab:SRMR compares the sound quality between prior CoNNear and dCoNNear for clean speech, noisy speech, and music, based on SRMR scores. For both clean and noisy speech, dCoNNear achieved higher SRMR scores than prior CoNNear, with the most significant improvement observed in noisy speech. This suggests that the dCoNNear-based model enhances the overall sound quality of the resulting speech signals. However, dCoNNear showed lower SRMR scores for music compared to prior CoNNear. This is likely due to the nature of music, which involves harmonics and overtones that are crucial for its perception. SRMR, however, primarily focuses on human vocal frequencies and modulations, and may not fully capture the complexities of music. Fig. 7 displays example spectrograms for each audio type across the different models. The prior CoNNear models exhibited tonal artifacts (horizontal lines) for all audio types, which were not present in the original inputs. These artifacts were most noticeable in clean speech. In contrast, the horizontal lines were significantly reduced in the spectrograms of the dCoNNear models. Although the SRMR scores indicated that the sound quality of music was better for prior CoNNear than for dCoNNear, the spectrograms of music revealed that dCoNNear introduced far fewer artifacts compared to prior CoNNear. This suggests that dCoNNear effectively reduces artifacts across different audio types. The corresponding audio samples of music, clean speech, and noisy speech are provided at the link.
TABLE III: Comparison of sound quality across different HA models based on SRMR.
Input dCoNNear prior CoNNear
Clean speech 7.39 6.77 6.13
Noisy speech 6.45 6.24 5.17
Music 1.93 1.75 1.91

### v3

@@ III-B Auditory modules and HA model
When selecting hyperparameters, two primary considerations should be addressed [baby2021convolutional, drakopoulos2023dnn]: the receptive field (RF) and the specific activation function. A larger receptive field (RF) ensures that the CNN-based network accurately simulates the auditory model’s impulse response. The selected activation functions need to capture the nonlinear characteristics of auditory processing, such as the compressive growth of the BM vibrations in the cochlear model and the rectification-like behavior of the IHCs. The framework’s hyperparameters are detailed in Table I. Specifically, $K_{1}$ and $K_{2}$ denote the kernel sizes of the dilated depthwise convolution layers (Dh-conv and Df-conv) within each memory block. The parameter $H$ indicates the number of channels in the hidden layers. Additionally, $L$, $L_{l}$, and $L_{r}$ represent the input length, left context, and right context, respectively.
The dCoNNearcochlear transforms acoustic signals into cochlear basilar-membrane vibrations. As described in [baby2021convolutional], the hyperbolic tangent (Tanh) activation function was used, because it allows the activations to resemble the input-output relations of the auditory model and supports both positive and negative deflections of the basilar membrane.
The dCoNNearIHC model emulates the IHCs’ function of sensing basilar membrane vibrations and converting them into IHC receptor potential changes. It employs a Tanh nonlinearity in the first stack and a sigmoid function in the second stack. This approach mirrors the compressive characteristics of the IHC input/output functions, with the sigmoid function effectively capturing the IHC receptor potential as a negative voltage difference [drakopoulos2023neural].
The dCoNNearANF model converts IHC receptor potentials into firing rates for three types of ANFs. The initial M blocks are shared layers, with each branch containing M blocks to independently generate HSR, MSR, and LSR fiber responses. The overall AN response ($r_{f}$) is derived by summing the three ANF responses. The ReLU activation function is applied here, as ANF processing differs from cochlear and IHC processing in that it only involves positive output values, without the compressive properties observed in the previous stages. The model features substantial left context windows of 7936 samples, ensuring accurate capture of the time-dependent recovery properties of the ANF models [drakopoulos2021convolutional].
@@
TABLE I: Hyperparamters of the dCoNNears and HA-model. M and R represent the number of memory blocks and the repetition number; K1 and K2 denote the kernel size of Dh-conv and Df-conv in each memory block; H indicates the channel number of hidden layers; Ll, and Lr represent the left context and right context, respectively
@@ III-C Individualisation of hearing impairment
To individualise hearing impaired pathway as shown in Fig. 2, we adjusted the normal hearing (NH) auditory modules to simulate different degrees of outer hair cell (OHC) loss and cochlear synaptopathy (CS) in the HI pathway. To simulate OHC loss, we retrained the NH cochlear model using transfer learning, based on a specific gain-loss profile or an individual audiogram [van2020hearing]. The NH auditory nerve (AN) response ($r_{f}$) is computed as a weighted sum of the three ANF responses ($H_{NH}$ = 13, $M_{NH}$ = 3, and $L_{NH}$ = 3). To simulate CS, the model was modified with adjusted weights $H_{HI}$, $M_{HI}$, and $L_{HI}$ [drakopoulos2023neural]. These personalized adjustments ensured the models accurately represented individual hearing impairments, providing tailored audio processing solutions that minimized artifacts and preserved fidelity.
@@ IV Training strategy
The training procedure involved two main stages: (1) training the auditory modules (dCoNNearcochlear, dCoNNearIHC, and dCoNNearANF), and (2) training the HA model based on the closed-loop framework including the NH and HI pathways. Firstly, we trained the auditory modules using simulated responses from analytical models as targets, following the procedure outlined in [baby2021convolutional, drakopoulos2021convolutional]. Once the auditory elements were trained, their parameters were frozen, and the HA models were then optimized within the closed-loop system as first described in [drakopoulos2023neural].
@@ IV-A Training auditory elements
The auditory module training targets were derived from various analytical models: a biophysical transmission line (TL) model for dCoNNearcochlear [verhulst2018computational], an analytical Hodgkin–Huxley-type model for dCoNNearIHC [altoe2017ihc], and a three-store diffusion model of the ANF synapse for dCoNNearANF [altoe2018ANF]. To train the models, we randomly selected 2,310 utterances from the full TIMIT speech corpus [garofolo1993TIMIT] without replacement. The resulting speaker distribution was approximately 70% male and 30% female, consistent with the overall gender distribution in TIMIT. The input signals were first upsampled to 100 kHz to accurately solve the reference models, and then downsampled to 20 kHz for training. The root-mean-square (RMS) energy of each sentence was adjusted to a specified sound pressure level (SPL) using the following equation:
@@ IV-A1 Training dCoNNearcochlear
The dCoNNearcochlear model was trained to simulate basilar membrane (BM) displacements for 201 center frequencies between 112 Hz and 12 kHz [verhulst2018computational], as human hearing sensitivity diminishes above 12 kHz. The RMS energy of each sentence was adjusted to $L_{D}=70$ dB SPL. The input signal and TL-model outputs were segmented into windows of 2,048 samples, with left ($L_{l}$) and right ($L_{r}$) contexts of 256 samples each, resulting in a total input length of 2,560 samples. To facilitate training, a scaling factor of $10^{6}$ was applied to the simulated BM displacements (expressed in $\mu m$]), ensuring that the datasets maintained a statistical mean near zero and a standard deviation close to one.
@@ IV-A2 Training dCoNNearIHC and dCoNNearANF
The training data for IHC and ANF, with dimension $N\textsubscript{CF}=201$, were converted into a one-dimensional dataset of 2,310 $\cdot N\textsubscript{CF}$ samples. This approach assumed that the reference IHC and ANF models had CF-independent parameters, while the BM displacements had CF-dependent impulse responses due to cochlear mechanics [drakopoulos2021convolutional].
@@ IV-C Loss function
The training strategies from [baby2021convolutional] and [drakopoulos2021convolutional] were employed to train the dCoNNearcochlear, dCoNNearIHC, and dCoNNearANF models. The loss function for these models was optimized using the mean absolute error (MAE) between the analytical outputs and the predicted dCoNNear outputs.
For training the HA model parameters, we utilized a combined loss function that accounted for different representations of the differences between the NH responses $r_{f}$ and the HI $\hat{r_{f}}$. As outlined in [drakopoulos2023neural], additional constraints, such as population responses, were introduced to the training process to minimize features functionally relevant to hearing loss and auditory perception, given the highly nonlinear nature of auditory processing. The combined loss function incorporated both the AN responses and the AN population responses across all simulated CFs:
@@ IV-D Hearing impaired elements
We optimized the HA model to compensate for a specific hearing-impaired (HI) profile “Slope35-7,0,0” [drakopoulos2023neural]. This profile represents a sloping high-frequency outer hair cell (OHC) loss beginning at 1 kHz, with hearing thresholds of 35 dB HL at 8 kHz, referred to as Slope35. Additionally, there is a complete loss of low spontaneous rate (LSR) and medium spontaneous rate (MSR) auditory nerve fibers (ANFs) (HHI = 0 and MHI= 0) and approximately 46% loss of high spontaneous rate (HSR) ANFs (LHI = 7) to simulate the effects of age-related cochlear synaptopathy.
@@ IV-E Training setup
The learning rates for model training was set to $10^{-4}$. If the validation loss did not decrease over five consecutive epochs, the learning rate was halved to ensure optimal training progression. All models were trained for 50 epochs using the Adam optimizer [kingma2014adam]. Training was conducted on three NVIDIA A30 GPUs using the PyTorch platform.
@@ V-A Evaluating Artifacts
To determine whether the artifacts caused by the prior CoNNear were eliminated, we analyzed 70-dB SPL 1-kHz tone responses at various auditory processing stages, as well as the HA models trained within the closed-loop system based on both prior CoNNear and dCoNNear architectures. We quantified artifacts using total harmonic distortion (THD) at each step of the closed-loop system [shmilovitz2005THD]. As aliased harmonics are generated by the prior CoNNear models in the 1-kHz tone responses as shown in Fig. 3, we examined harmonics at one-quarter intervals of the fundamental frequency:
@@
The Speech-to-Reverberation Modulation Energy Ratio (SRMR) [falk2010SRMR], DNSMOS (overall) [reddy2021dnsmos] and Perceptual evaluation of speech quality (PESQ) [pesq2001perceptual] were used to evaluate the sound quality of the resulting HA models trained using both the prior CoNNear and dCoNNear-based frameworks. For all metrics, higher scores indicate higher sound quality. We evaluated sound quality across three audio types: clean speech, noisy speech, and music. The clean speech dataset consisted of 550 utterances from TIMIT and LibriTTS [zen2019libritts], which were independent of the training set of HA models. The one-hour noisy speech dataset was generated by mixing the clean speech samples and the noise from Freesound [fonseca2017freesound] at SNR of 0 dB. We used environmental noise from the Freesound database, which typically includes sounds such as street noise, sounds of people talking, machinery hum, and ambient indoor sounds. The music dataset consisted of 100 randomly selected 30-second tracks from FMA [fma_dataset].
@@ V-B Evaluating Auditory Elements
For the dCoNNearcochlear, we simulated the QERB and basilar membrane (BM) excitation patterns, as described in [baby2021convolutional]. The QERB metric characterizes the level-dependent cochlear filter properties derived from the BM impulse response and is calculated as follows:
@@
For the dCoNNearANF, rate-level curves and synchrony levels were measured [drakopoulos2023neural]. Rate-level curves evaluate ANF responses to changes in stimulus level. We generated pure-tone stimuli (50-ms duration, 2.5-ms rise/fall ramp) with levels between 0 and 100 dB at frequencies of approximately 1 and 4 kHz, based on the corresponding CFs of the ANF models (1007 and 3972.7 Hz). The rate-level functions were derived by computing the average response 10–40 ms after stimulus onset. The ANF synchrony level describes the non-monotonic relation between ANF response and the stimulus level. Fully modulated 400-ms long pure tones with a modulation frequency $f_{m}$ of 100 Hz and carrier frequencies of 3972.7 Hz (henceforth referred to as 4 kHz) were simulated, and the synchrony-level functions were calculated by extracting the magnitude of the $f_{m}$ component from the Fourier spectrum of the fibers’ firing rate.
@@ V-C Evaluating HA models
To evaluate the restoration performance of the trained HA models, we used normalized root-mean-square error (NRMSE) in [drakopoulos2023neural], which was computed between the simulated NH and HI AN population responses, normalised to the maximum of the NH response for each sentence:
@@
NRMSE calculations were performed on 550 randomly selected clean speech utterances from the TIMIT and LibriTTS [zen2019libritts] utterances, which were excluded from the HA training dataset. These utterances had RMS energy levels ranging from 40 to 70 dB SPL, in 10 dB increments.
@@ VI Results
In this section, we evaluated the proposed artifact-free dCoNNear-based closed-loop framework, designed for individualized DNN-based hearing-aid (HA) models. The evaluation aimed to determine whether the artifacts associated with the prior CoNNear architecture were successfully eliminated in the dCoNNear-based auditory elements. We also compared the sound quality of the HA models trained between the prior CoNNear and new dCoNNear frameworks. Additionally, we assessed the biophysical properties of the auditory elements using stimuli that were not part of the training material. Finally, we compared the hearing-aid restoration performance of the models developed using the prior CoNNear-based framework with those developed using the dCoNNear framework.
@@
Figure 8: Comparison of the frequency responses of a 1-kHz pure tone by different models.
@@ VI-A Artifacts and sound quality
Table III and IV compare the sound quality between prior CoNNear and dCoNNear for clean speech, noisy speech, and music, using two different corpora: TIMIT and LibriTTS. Across both datasets, dCoNNear consistently outperformed prior CoNNear in terms of SRMR, DNSMOS, and PESQ. This consistent improvement in both clean and noisy conditions indicates that the dCoNNear-based model enhances the overall sound quality of the resulting speech signals. For music, however, the trend deviated slightly. While dCoNNear still outperformed prior CoNNear in DNSMOS and PESQ, its SRMR scores were marginally lower. This is likely due to the nature of music, which involves harmonics and overtones that are crucial for its perception. SRMR, however, primarily focuses on human vocal frequencies and modulations, and may not fully capture the complexities of music.
Fig. 7 presents spectrogram examples across different audio types for both model architectures. The prior CoNNear models exhibited prominent tonal artifacts, visible as horizontal lines over time, which were absent in the original inputs. These artifacts were especially pronounced in clean speech and could be attributed to two primary structural issues: (1) tonal artifacts introduced by transposed convolution layers and (2) spectral replication, commonly known as imaging artifacts, caused by naive upsampling operations. In the spectrograms of speech processed by the prior CoNNear-based model, equally spaced mirror-image bands appeared in the high-frequency regions (highlighted by the white rectangular box). These image artifacts resulted from spectral replication during upsampling. In contrast, the horizontal lines and replicated high-frequency bands were minimized in the dCoNNear outputs, indicating that the architecture effectively suppressed these types of artifacts.
Fig. 8 shows the frequency responses for a 1 kHz pure tone input processed by both models. The prior CoNNear output (blue) exhibited dense clusters of spurious low-frequency peaks below 500 Hz, which were absent in both the input and the dCoNNear output. These peaks reflected aliasing artifacts caused by repeated downsampling without adequate anti-aliasing filtering. Specifically, the prior CoNNear-based HA model includes eight downsampling layers with stride 2, reducing the effective Nyquist frequency at the deepest layer to approximately 39 Hz. As a result, the original 1 kHz tone was repeatedly folded into lower frequency bands at each stage, accumulating into artificial low-frequency components. These artifacts could perceptually manifest as low-frequency buzzing or modulation noise. In contrast, dCoNNear avoids the aliasing artifacts due to the problematic downsampling, yielding a cleaner and perceptually more natural output.
The resulting audio samples of music, clean speech, and noisy speech are provided online  [^1] https://github.com/chuan997/Trans24_audio_samples.
TABLE III: Comparison of sound quality across different HA models where the clean speech corpus is from TIMIT dataset.
Aduio type Model SRMR DNSMOS PESQ
Clean speech Unprocessed 7.39 3.33 4.64
prior CoNNear 6.13 2.22 1.82
dCoNNear 6.77 3.26 4.31
Noisy speech Unprocessed 6.45 2.64 2.68
prior CoNNear 5.17 2.06 2.34
dCoNNear 6.24 2.52 2.62
Music Unprocessed 1.93 1.09 -
prior CoNNear 1.91 1.08 1.75
dCoNNear 1.75 1.07 2.25
TABLE IV: Comparison of sound quality across different HA models where the clean speech corpus is from LibriTTS dataset.
Aduio type Model SRMR DNSMOS PESQ
Clean speech Unprocessed 7.84 3.49 4.64
prior CoNNear 6.48 2.24 1.97
dCoNNear 7.01 3.18 4.22
Noisy speech Unprocessed 6.71 2.57 2.73
prior CoNNear 5.46 2.11 2.24
dCoNNear 6.41 2.49 2.61